# CroMe: Multimodal Fake News Detection using Cross-Modal Tri-Transformer and Metric Learning

### v1

@@ 3.4 Cross-Modal Tri-Transformer Fusion
Cosine Similarity is computed between different combinations of the inputs $Z_{\text{i2}}$, $Z_{\text{t2}}$, and $Z_{\text{b}}$ to measure the relationships between the modalities. The similarities $\mathcal{S}_{\text{ti}}$ ($Z_{\text{t2}}$, $Z_{\text{i2}}$), $\mathcal{S}_{\text{tb}}$ ($Z_{\text{t2}}$, $Z_{\text{b}}$), and $\mathcal{S}_{\text{ib}}$ ($Z_{\text{i2}}$, $Z_{\text{b}}$) are computed using cosine similarity, defined for two vectors $\mathbf{a}$ and $\mathbf{b}$ as:
$\mathcal{S}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\times\|\mathbf{b}\|}\,.$
The combined similarity is computed by weighting individual similarity scores with learnable parameters $w_{\text{i2}}$, $w_{\text{b}}$, $w_{\text{t2}}$, and adding a bias term $b$ to account for modality differences. This similarity is processed through linear layers, ReLU activations, and batch normalization, producing $\mathbf{C}_{\text{2}}$. BLIP2 components use cosine similarity as it effectively captures directional alignment, simplifies the fusion process, and maintains key correlations between modalities Li et al. (2023). The outputs $\mathbf{C}_{\text{1}}$ and $\mathbf{C}_{\text{2}}$ are combined into $Z_{\text{c}}$ then processed through fully connected layers with batch normalization, ReLU activations, and dropout, resulting in the final correlation output used in the subsequent Tri-Transformer.
Tri-Transformer
The text and image embeddings ($Z_{\text{i}}$ and $Z_{\text{t}}$) are passed through their respective projection networks and, along with correlation outputs ($Z_{\text{c}}$), are fed into the Tri-Transformer (Figure 4(c)). This framework is based on the Multi-Modal Tri-Transformer in Choi and Kim (2024). Unlike TT-BLIP’s text-focused cross-modal attention, this method processes text, image, and correlation independently.
$f_{*}=\text{MultiHead}\left(\text{Softmax}\left(\frac{Q_{*}K_{*}^{\mathbf{T}}}{\sqrt{d_{h}}}\right)\times V_{*};\theta_{\text{att}}^{*}\right),$
where $Q_{*},K_{*},V_{*}$ are query, key, and value matrices, and $\theta_{\text{att}}^{*}$ are attention parameters. The outputs processed through Multi-Layer Perceptron (MLP) layers and combined into a unified tensor.
Table 1: Experimental Results for Weibo, Weibo-21 and Politifact Datasets. A dash (‘-’) signifies that results are not provided in their corresponding research paper.
Datasets Methods Accuracy Fake News Real News
@@
Weibo EANN Wang et al. (2018) 0.827 0.847 0.812 0.829 0.807 0.843 0.825
MVAE Khattar et al. (2019) 0.824 0.854 0.769 0.809 0.802 0.875 0.837
MPFN Jing et al. (2023) 0.838 0.857 0.894 0.889 0.873 0.863 0.876
Spotfake Singhal et al. (2019) 0.892 0.902 0.964 0.932 0.847 0.656 0.739
SAFE Zhou et al. (2020) 0.762 0.831 0.724 0.774 0.695 0.811 0.748
BDANN Zhang et al. (2020) 0.821 0.790 0.610 0.690 0.830 0.920 0.870
LIIMR Singhal et al. (2022) 0.900 0.882 0.823 0.847 0.908 0.941 0.925
MCAN Wu et al. (2021) 0.899 0.913 0.889 0.901 0.884 0.909 0.897
CAFE Chen et al. (2022) 0.840 0.855 0.830 0.842 0.825 0.851 0.837
FND-CLIP Zhou et al. (2023) 0.907 0.914 0.901 0.908 0.914 0.901 0.907
TT-BLIP Choi and Kim (2024) 0.961 0.979 0.944 0.961 0.944 0.980 0.962
@@
Weibo-21 HAN Yang et al. (2016) 0.792 - - 0.792 - - -
Vaibhav(2019) Vaibhav et al. (2019) 0.771 - - 0.771 - - -
FakeFlow Ghanem et al. (2021) 0.781 - - 0.780 - - -
EANN Wang et al. (2018) 0.870 0.902 0.825 0.862 0.841 0.912 0.875
SpotFake Singhal et al. (2019) 0.851 0.953 0.733 0.828 0.786 0.964 0.866
CAFE Chen et al. (2022) 0.882 0.857 0.915 0.885 0.907 0.844 0.876
CroMe (Ours) 0.917 0.944 0.917 0.930 0.880 0.918 0.930
Politifact RoBERTa-MWSS Shu et al. (2004) 0.820 - - 0.820 - - -
SAFE Zhou et al. (2020) 0.874 0.851 0.830 0.840 0.889 0.903 0.896
Spotfake+ Singhal et al. (2020) 0.846 - - - - - -
TM Bhattarai et al. (2022) 0.871 - - - 0.901 - -
LSTM-ATT Lin et al. (2019) 0.832 0.828 0.832 0.830 0.836 0.832 0.829
DistilBert Allein et al. (2021) 0.741 0.875 0.636 0.737 0.647 0.880 0.746
CAFE Chen et al. (2022) 0.864 0.724 0.778 0.750 0.895 0.919 0.907
FND-CLIP Zhou et al. (2023) 0.942 0.897 0.897 0.897 0.960 0.960 0.960
TT-BLIP Choi and Kim (2024) 0.904 0.913 0.724 0.808 0.901 0.973 0.936
@@
3.5 Fake News Detector
The fake news detector uses the fusion representation to predict whether news articles are fake or real. This fused representation is processed by a classifier composed of three fully connected layers, ReLU activations, and batch normalization, resulting in a binary classification. The classifier assigns labels of ‘real’ (0) or ‘fake’ (1) to the news content, where $\hat{y}=[\hat{y}_{0},\hat{y}_{1}]$ denotes the probabilities of the news being ‘real’ (0) or ‘fake’ (1). Cross-entropy is used to define the loss function $L(\theta)$, where $\theta$ represents the model parameters.
$L(\theta)=-y\log(\hat{y}_{1})-(1-y)\log(\hat{y}_{0})$
The total loss $L_{\text{total}}$ is defined, where $\alpha$ controls the weight of this metric loss. The optimal value of $\beta$ determined through a grid search over the range [0.1, 1] using a step 0.1. The model aims to minimize the total loss $L_{\text{total}}$ for each news data by learning $\theta$ through back propagation.
$L_{\text{total}}=L(\theta)+\beta\cdot L(X)$
4 Experiments
4.1 Experimental Setup
Datasets
To evaluate the performance of CroMe, three datasets are used: Weibo Jin et al. (2017), Weibo-21 Nan et al. (2021), and Politifact Shu et al. (2020). The Weibo dataset contains 6,137 training articles (2,802 fake and 3,335 real) and 1,685 test articles (833 fake and 852 real). The Weibo-21 dataset includes 4,640 real and 4,487 fake articles, split into training and testing sets with an 8:2 ratio. The Politifact dataset consists of 381 training articles (246 fake and 135 real) and 104 test articles (74 fake and 30 real).
Training Settings
Text encoding utilized the pretrained BERT model Devlin et al. (2019) for Chinese in the Weibo and Weibo-21 datasets and the ”bert-base-uncased” model for the Politifact dataset. Images were resized to $224\times 224$ pixels and encoded using Masked Autoencoders (MAE) He et al. (2022). Image-text pairs were encoded with the pretrained BLIP2 model Li et al. (2023), translating Chinese texts to English via the Google Translation API Johnson (2012). The model employed five Adam optimizers for $Z_{\text{*}}\in\{Z_{\text{i1}},Z_{\text{i2}},Z_{\text{t1}},Z_{\text{t2}},Z_{\text{b}}\}$, with a learning rate of $1\times 10^{-3}$, a batch size of 64, and was trained for 50 epochs. Metric Learning parameters were set as follows: an iterative frequency of 5 epochs per modality, $\alpha=16$, $\delta=0.1$, and $\beta=0.1$, balancing classification and metric losses. These values were determined through preliminary experiments to optimize performance.
4.2 Results and Analysis
CroMe’s performance is evaluated against state-of-the-art models shown in Table 1. The evaluation metrics used include accuracy and precision, recall, and F1 scores for both real and fake news. CroMe achieved the highest accuracy of 0.974 in the experiments using Weibo, surpassing TT-BLIP by 1.3% and FND-CLIP by 6.7%. Similarly, when using Weibo-21, CroMe reached an accuracy of 0.917, outperforming CAFE by 3.5% and EANN by 4.7%. For both datasets, CroMe ranked 1st or 2nd in precision, recall, and F1 scores for both fake and real news. CroMe achieves 0.933 accuracy that closely matches FND-CLIP’s 0.942. The reason may be that the dataset size is too small.
@@ 4.2 Results and Analysis
Table 2: Ablation experimental results of CroMe.
@@
MAE BLIP2Img BERT BLIP2Txt BLIP2Img,Txt CM MT TT Fake News Real News
@@
CroMe excelled in fake news detection due to three key factors: 1) Advanced multimodal feature extraction through the BLIP2 model enhances CroMe’s capabilities by capturing detailed features from both text and images, allowing for more precise feature extraction. This precision improves the model’s ability to distinguish between real and fake news. 2) The Cross-Modal Tri-Transformer Fusion (CMTTF) computes and integrates cross-modal similarities, enhancing the model’s ability to utilize interactions between different modalities which is crucial for accurate fake news detection. 3) CroMe incorporates a Metric Learning module using the Proxy Anchor method, which focuses on intra-modality relationships. This ensures that features within the same modality are more closely aligned, leading to better representation learning and overall improved performance in detecting fake news. For smaller datasets, future research could explore data augmentation or transfer learning to improve performance.
4.2.1 Proxy Anchor Loss Parameter Analysis
The analysis of the Weibo, Weibo-21, and Politifact datasets under different Proxy Anchor Loss hyperparameters, $\delta$ (margin) and $\alpha$ (alpha), is shown in Figure 5. As the margin increases, performance declines, with larger margins causing greater drops. Performance improves as alpha approaches 16, with the best results at alpha 16. Performance is more sensitive to margin variations than alpha. Alpha was tested at 4, 8, 16, 32, and margin at 0.1, 0.2, 0.3, 0.4. The best performance was achieved with alpha 16 and margin 0.1.
4.3 Ablation
@@
    Without Image (BLIP2Img + MAE): Removing the image encoders and using only text encoders and other components.
@@
    Without Text (BLIP2Txt + BERT): Removing the text encoders and using only image encoders and other components.
@@
    Without BLIP2 (BLIP2Img + BLIP2Txt + BLIP2Img-Txt): Removing all BLIP2 components (text, image and image-text) and using only other encoders.
@@
    Without BLIP2Img-Txt: Removing the image-text feature component of BLIP2, retaining the individual text and image BLIP2 encoders.
@@
The ablation study highlights three components that most significantly impact accuracy when removed. First, excluding the BLIP2 text components causes a substantial drop in accuracy across all datasets, as they are critical for understanding textual context necessary for identifying fake news. Second, removing the Cross-Modal Fusion (CM) component weakens performance by preventing effective integration of inter-modal fusion and cosine similarity features from text and images. Lastly, the Metric Learning module (MT) is essential for learning distinct features that differentiate fake and real news; its removal reduces accuracy. Overall, uni-modal and cross-modal features, along with CMTTF and Metric Learning, are crucial for the model’s effectiveness in detecting fake news.
4.3.1 T-SNE visualizations
@@
Figure 6 provides a T-SNE Van der Maaten and Hinton (2008) visualization of features before classification, comparing various CroMe settings: CroMe w/o image, CroMe w/o text, CroMe w/o BLIP, CroMe w/o fusion, CroMe w/o metric and the full CroMe model on the Weibo test dataset. Dots of the same color indicate the same label. Figure 6(d) shows that without the Cross-Modal Tri-Transformer Fusion (CMTTF) module, fake and real news instances are not well-separated. Including it improves clustering and distinction between classes. Comparisons of Figures 6(a)6(b)6(c), and 6(f) shows that removing BLIP2 feature extraction results in less clear clustering, highlighting its importance for integrating image, text, and image-text data. Excluding text features also reduces separation in t-SNE plots, showing that image and image-text features alone are less effective. Comparing CroMe w/o MT (Metric Learning, Figure 6(e)) and CroMe (Figure 6(f)), the degree of separation of the sample dots in Figure 6(f) is higher. This indicates that capturing intra-modality relationships ensures effective representation learning within the same modality.
5 Conclusion

### v2

@@ III-D1 Cross-Modal Fusion
Second, Cosine Similarity is computed between different combinations of the inputs $Z_{\text{i2}}$, $Z_{\text{t2}}$, and $Z_{\text{b}}$ to measure the relationships between the modalities. The similarities $\mathcal{S}_{\text{ti}}$ ($Z_{\text{t2}}$, $Z_{\text{i2}}$), $\mathcal{S}_{\text{tb}}$ ($Z_{\text{t2}}$, $Z_{\text{b}}$), and $\mathcal{S}_{\text{ib}}$ ($Z_{\text{i2}}$, $Z_{\text{b}}$) are computed using cosine similarity, defined for two vectors $\mathbf{a}$ and $\mathbf{b}$ as:
$\mathcal{S}(\mathbf{a},\mathbf{b})=\frac{\mathbf{a}\cdot\mathbf{b}}{\|\mathbf{a}\|\times\|\mathbf{b}\|}\,.$ (5)
The combined similarity is computed by weighting individual similarity scores with learnable parameters $w_{\text{i2}}$, $w_{\text{b}}$, $w_{\text{t2}}$, and adding a bias term $b$ to account for modality differences. This similarity is processed through linear layers, ReLU activations, and batch normalization, producing $\mathbf{C}_{\text{2}}$. BLIP2-OPT components use cosine similarity as it effectively captures directional alignment, simplifies the fusion process, and maintains key correlations between modalities [li2023blip]. The outputs $\mathbf{C}_{\text{1}}$ and $\mathbf{C}_{\text{2}}$ are combined into $Z_{\text{c}}$, then processed through fully connected layers with batch normalization, ReLU activations, and dropout, resulting in the final correlation output used in the subsequent Tri-Transformer.
III-D2 Tri-Transformer
The text and image embeddings ($Z_{\text{i}}$ and $Z_{\text{t}}$) are passed through their respective projection networks and, along with correlation outputs ($Z_{\text{c}}$), are fed into the Tri-Transformer (as shown in Figure 4(c)). This framework is based on the Multimodal Tri-Transformer as discussed in [10706486]. Unlike TT-BLIP’s text-focused cross-modal attention, this method processes text, image, and correlation independently.
$f_{*}=\text{MultiHead}\left(\text{Softmax}\left(\frac{Q_{*}K_{*}^{\mathbf{T}}}{\sqrt{d_{h}}}\right)\times V_{*};\theta_{\text{att}}^{*}\right)$ (6)
Here, $Q_{*},K_{*},V_{*}$ are query, key, and value matrices, $\theta_{\text{att}}^{*}$ are attention parameters, and $\in\{t,i,c\}$ corresponds to text, image, and correlation outputs respectively. Plus, $d_{h}$ represents the dimensionality, with $\mathbf{T}$ indicating the transpose operation. The outputs processed through Multi-Layer Perceptron (MLP) layers and combined into a unified tensor.
TABLE I: Experimental results for Weibo, Weibo-21 and Politifact Datasets. A dash (‘-’) signifies that results are not provided in their corresponding research paper.
Datasets Models Accuracy Fake News Real News
@@ III-D2 Tri-Transformer
Weibo EANN [wang2018eann] 0.827 0.847 0.812 0.829 0.807 0.843 0.825
MVAE [khattar2019mvae] 0.824 0.854 0.769 0.809 0.802 0.875 0.837
MPFN [jing2023multimodal] 0.838 0.857 0.894 0.889 0.873 0.863 0.876
Spotfake [singhal2019spotfake] 0.892 0.902 0.964 0.932 0.847 0.656 0.739
SAFE [zhou2020similarity] 0.762 0.831 0.724 0.774 0.695 0.811 0.748
BDANN [zhang2020bdann] 0.821 0.790 0.610 0.690 0.830 0.920 0.870
LIIMR [singhal2022leveraging] 0.900 0.882 0.823 0.847 0.908 0.941 0.925
MCAN [wu2021multimodal] 0.899 0.913 0.889 0.901 0.884 0.909 0.897
CAFE [chen2022cross] 0.840 0.855 0.830 0.842 0.825 0.851 0.837
FND-CLIP [zhou2023multimodal] 0.907 0.914 0.901 0.908 0.914 0.901 0.907
TT-BLIP [10706486] 0.961 0.979 0.944 0.961 0.944 0.980 0.962
@@
Weibo-21 EANN [wang2018eann] 0.870 0.902 0.825 0.862 0.841 0.912 0.875
SpotFake [singhal2019spotfake] 0.851 0.953 0.733 0.828 0.786 0.964 0.866
CAFE [chen2022cross] 0.882 0.857 0.915 0.885 0.907 0.844 0.876
CroMe (Ours) 0.917 0.944 0.917 0.930 0.880 0.918 0.899
Politifact RoBERTa-MWSS [shu2004leveraging] 0.820 - - 0.820 - - -
SAFE [zhou2020similarity] 0.874 0.851 0.830 0.840 0.889 0.903 0.896
Spotfake+ [singhal2020spotfake+] 0.846 - - - - - -
TM [bhattarai-etal-2022-explainable] 0.871 - - - 0.901 - -
LSTM-ATT [lin2019detecting] 0.832 0.828 0.832 0.830 0.836 0.832 0.829
DistilBert [allein2021like] 0.741 0.875 0.636 0.737 0.647 0.880 0.746
CAFE [chen2022cross] 0.864 0.724 0.778 0.750 0.895 0.919 0.907
FND-CLIP [zhou2023multimodal] 0.942 0.897 0.897 0.897 0.960 0.960 0.960
TT-BLIP [10706486] 0.904 0.913 0.724 0.808 0.901 0.973 0.936
@@
III-E Fake News Detector
Fake news detector uses the fusion representation to predict whether news articles are fake or real. This fused representation is processed by a classifier composed of three fully connected layers, ReLU activations, and batch normalization, resulting in a binary classification. The classifier assigns labels of ‘real’ (0) or ‘fake’ (1) to the news content, where $\hat{y}=[\hat{y}_{0},\hat{y}_{1}]$ denotes the probabilities of the news being ‘real’ or ‘fake’. Cross-entropy is used to define the loss function $L(\theta)$, where $\theta$ represents the model parameters, as follows,
$L(\theta)=-y\log(\hat{y}_{1})-(1-y)\log(\hat{y}_{0})\,.$ (7)
The total loss $L_{\text{total}}$ is defined, where $\beta$ controls the weight of this metric loss (as described in Equation 1). The optimal value of $\beta$ was determined through a grid search over the range [0.1, 1] using a step of 0.1. The model aims to minimize the total loss $L_{\text{total}}$ for each news data by learning $\theta$ through back-propagation.
$L_{\text{total}}=L(\theta)+\beta\cdot L(X)$ (8)
IV Evaluation
IV-A Experimental Setup
IV-A1 Datasets
To evaluate the performance of CroMe, three datasets are used: Weibo [jin2017multimodal], Weibo-21 [nan2021mdfend], and Politifact [shu2020fakenewsnet]. The Weibo dataset contains 6,137 training articles (2,802 fake and 3,335 real) and 1,685 test articles (833 fake and 852 real) [jin2017multimodal]. The Weibo-21 dataset includes 4,640 real and 4,487 fake articles, split into training and testing sets with an 8:2 ratio [nan2021mdfend]. The Politifact dataset consists of 381 training articles (246 fake and 135 real) and 104 test articles (74 fake and 30 real) [shu2020fakenewsnet]. Each dataset (Weibo, Weibo-21, and Politifact) was used separately for training and evaluation to avoid any data overlap or leakage.
IV-A2 Training Settings
Text encoding utilized the pretrained BERT model [devlin-etal-2019-bert] for Chinese in the Weibo and Weibo-21 datasets and the ”bert-base-uncased” model for the Politifact dataset. Images were resized to $224\times 224$ pixels and encoded using Masked Autoencoders (MAE) [he2022masked]. Image-text pairs were encoded with the pretrained BLIP2-OPT model [li2023blip], translating Chinese texts to English via the Google Translation API [doi:10.1080/07317131.2012.650971]. The model employed five Adam optimizers for $Z_{\text{*}}\in\{Z_{\text{i1}},Z_{\text{i2}},Z_{\text{t1}},Z_{\text{t2}},Z_{\text{b}}\}$, with a learning rate of $1\times 10^{-3}$, a batch size of 64, and was trained for 50 epochs. Metric Learning parameters were set as follows: an iterative frequency of 5 epochs per modality, $\alpha=16$, $\delta=0.1$, and $\beta=0.1$, balancing classification and metric losses. These values were determined through preliminary experiments to optimize performance. Experiments were performed five times, and the results are averaged across these runs.
IV-B Results and Analysis
CroMe’s performance is evaluated against state-of-the-art models shown in Table I. CroMe results were obtained from our implementation, while baseline results were cited from their respective original papers using the same datasets and evaluation metrics. The evaluation metrics used include accuracy and precision, recall, and F1 scores for both real and fake news. CroMe achieved the highest accuracy of 0.974 in the experiments using Weibo, surpassing TT-BLIP by 1.3% and FND-CLIP by 6.7%. Similarly, when using Weibo-21, CroMe reached an accuracy of 0.917, outperforming CAFE by 3.5% and EANN by 4.7%. For both datasets, CroMe ranked 1st or 2nd in precision, recall, and F1 scores for both fake and real news. CroMe achieves 0.933 accuracy that closely matches FND-CLIP’s 0.942. The reason may be that the dataset size is too small.
TABLE II: Ablation experimental results of CroMe.
@@ IV-B Results and Analysis
MAE BLIP2-OPTimg BERT BLIP2-OPTtxt BLIP2-OPTimg,txt CM MT $T$ Fake News Real News
@@
CroMe excelled in fake news detection due to three key factors: 1) Advanced multimodal feature extraction through the BLIP2-OPT model enhances CroMe’s capabilities by capturing detailed features from both text and images, allowing for more precise feature extraction. This precision improves the model’s ability to distinguish between real and fake news. 2) The Cross-Modal Tri-Transformer Fusion (CMTTF) computes and integrates cross-modal similarities, enhancing the model’s ability to utilize interactions between different modalities which is crucial for accurate fake news detection. 3) CroMe incorporates a Metric Learning module using the Proxy Anchor method, which focuses on intra-modality relationships. This ensures that features within the same modality are more closely aligned, leading to better representation learning and overall improved performance in detecting fake news. For smaller datasets, future research could explore data augmentation or transfer learning to improve performance.
The analysis of Weibo, Weibo-21, and Politifact datasets under different Proxy Anchor Loss hyperparameters, $\delta$ (margin) and $\alpha$ (alpha), is shown in Figure 5. As the margin increases, performance declines, with larger margins causing greater drops. Performance improves as alpha approaches 16, with the best results at alpha 16. Performance is more sensitive to margin variations than alpha. Alpha was tested at 4, 8, 16, 32, and margin at 0.1, 0.2, 0.3, 0.4. The best performance was achieved with alpha 16 and margin 0.1.
IV-C Ablation
@@
    Without Image (BLIP2-OPTimg + MAE): Removing the image encoders and using only text encoders and other components.
@@
    Without Text (BLIP2-OPTtxt + BERT): Removing the text encoders and using only image encoders and other components.
@@
    Without BLIP2-OPT (BLIP2-OPTimg + BLIP2-OPTtxt + BLIP2-OPTimg,txt): Removing all BLIP2-OPT components (text, image and image-text) and using only other encoders.
@@
    Without BLIP2-OPTimg,txt: Removing the image-text feature component of BLIP2-OPT, retaining the individual text and image BLIP2-OPT encoders.
@@
The ablation study highlights three components that most significantly impact accuracy when removed. First, excluding the BLIP2-OPT text components causes a substantial drop in accuracy across all datasets, as they are critical for understanding textual context necessary for identifying fake news. Second, removing the Cross-Modal Fusion (CM) component weakens performance by preventing effective integration of inter-modal fusion and cosine similarity features from text and images. Lastly, the Metric Learning module (MT) is essential for learning distinct features that differentiate fake and real news; its removal reduces accuracy. The ablation settings (Without Image, Without Text and Without CM) correspond respectively to fake news cases involving image tampering, text manipulation, and image–text mismatch, validating the capability of CroMe’s cross-modal fusion method. While the Tri-Transformer (TT) contributes to overall performance, the CM and MT modules distinguish CroMe from prior models such as TT-BLIP. CM improves cross-modal fusion, and MT enhances feature clustering within each modality. These modules result in improved classification performance, contributing to higher multimodal classification accuracy.
@@
Figure 6 provides a T-SNE [van2008visualizing] visualization of features before classification, comparing various CroMe settings: CroMe w/o image, CroMe w/o text, CroMe w/o BLIP, CroMe w/o fusion, CroMe w/o metric and the full CroMe model on the Weibo test dataset. Dots of the same color indicate the same label. Figure 6(d) shows that without the Cross-Modal Tri-Transformer Fusion (CMTTF) module, fake and real news instances are not well-separated. Including it improves clustering and distinction between classes. Comparisons of Figures 6(a)6(b)6(c), and 6(f) shows that removing BLIP2-OPT feature extraction results in less clear clustering, highlighting its importance for integrating image, text, and image-text data. Excluding text features also reduces separation in t-SNE plots, showing that image and image-text features alone are less effective. Comparing CroMe w/o MT (Metric Learning, Figure 6(e)) and CroMe (Figure 6(f)), the degree of separation of the sample dots in Figure 6(f) is higher. This indicates that capturing intra-modality relationships ensures effective representation learning within the same modality.
IV-D Discussion
CroMe Architecture relies on machine translation when encoding Chinese posts using BLIP2-OPT. This dependency can introduce translation artifacts and culture-specific semantic drift. In addition, the Politifact dataset split (381/104 train/test) is comparatively small, and our slightly lower accuracy on this dataset suggests sensitivity to limited data. Future work should validate CroMe on larger and more diverse corpora, include cross-dataset transfer tests, and assess robustness when modalities are missing at inference (text-only or image-only). Data augmentation or transfer learning may mitigate the small-data effect.
V Conclusion